OpenStudy (anonymous):

Triangle ABC is congruent to triangle DEF. In triangle ABC, side AB measures 3, side BC measures x+13, and side CA measures 6. In triangle DEF, side DE measures 3, side EF measures 3x-21, and side FD measures 6. What equation would help you to solve for the side length of BC and EF? Explain your reasoning using complete sentences. (10 points)

OpenStudy (anonymous):

Im stuck. D:

OpenStudy (anonymous):

 so we take ef and bc x+13=3x-21 move xs to one side x-x+13=3x-x-21 13=2x-21 add 21 13+21=2x-21+21 x=24

OpenStudy (anonymous):

2x=24* x=12

OpenStudy (anonymous):

OH OKAY. THAT MAKES SENSE. wow. Thankyou so much. :) ill let you know if i have another one.

OpenStudy (anonymous):

:D Ill help anytime

OpenStudy (anonymous):

Isosceles trapezoid ABCD is constructed with line ZX drawn through its center. If the isosceles trapezoid is dilated using a scale factor of 1/3 and a line is drawn through the center of the new dilated figure, what relationship will the new line have with line ZX? Explain your reasoning using complete sentences. So i'm thinking it would stay the same. But. i'm not 100%.

OpenStudy (anonymous):

^^^ @BreBanana

OpenStudy (anonymous):

The line would stay the same if it is a continuous line, but if it was a line segment it would also shrink

OpenStudy (anonymous):

So I should just put.. Assuming that it's a continuous line, it would stay the same. But if it's a line segment then it would also dilate with the figure.

OpenStudy (anonymous):

That's what I would put.